# A Bayesian hierarchical model for methane emission source apportionment

### v1

@@ 3.1.3 Parameter layer
$\theta_{i}$ $\sim\text{Beta}(a_{i},b_{i})$
[…]
$r$ $\sim\text{Uniform}(0,1).$
Fixed hyperparameters $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$ control our prior belief about the probability that $\beta_{i}$ is non-zero and about the scale of $\beta_{i}$. For example, heaters on oil and gas sites are not used in the summer, and as such the $a_{i}$ and $b_{i}$ hyperparameters could be set such that $\theta_{i}$ has more mass close to zero in the prior during these months. By default, however, we use a uniform prior for $\theta_{i}$ by setting $a_{i}=b_{i}=1$ and relatively uninformative priors for the remaining parameters by setting $c_{i}=d_{i}=\alpha_{1}=\alpha_{2}=1$. The prior for $\sigma^{2}$ was set such that, in the prior model, the $\boldsymbol{\epsilon}$ are distributed according to a Student’s t-distribution with $\nu$ degrees of freedom, as we discovered that the $\boldsymbol{\epsilon}$ can be heavier tailed than a Gaussian through basic model diagnostics.
@@ 3.2 Posterior inference
For methane emission source apportionment on oil and gas sites, the joint posterior distribution for the emission rates, $\boldsymbol{\beta}$, and the indicator variables, $\boldsymbol{z}$, are the most useful model output. If we let $\boldsymbol{\xi}$ be a vector of all model parameters, $\boldsymbol{\xi}=\{\beta_{1},...,\beta_{p},z_{1},...,z_{p},\theta_{1},...,\theta_{p},\tau_{1}^{2},...,\tau_{p}^{2},\sigma^{2},\nu,r\}$, then Bayes’ theorem gives us the full joint distribution: $p(\boldsymbol{\xi}|\boldsymbol{y})\propto p(\boldsymbol{y}|\boldsymbol{\xi})p(\boldsymbol{\xi})$, where $p(\cdot)$ denotes a probability density. We obtain samples from the joint posterior using a Markov chain Monte Carlo (MCMC) algorithm. We describe our MCMC algorithm below, but see Gelman et al. (2015) for a more complete discussion of Bayesian inference via MCMC.
We sample from the joint posterior using a Metropolis-within-Gibbs algorithm. We are able to derive a closed form expression for the conditional distributions of all model parameters except $\nu$ and $r$, resulting in a fast sampler that mixes well. For $\nu$ and $r$, we update using two Metropolis-Hastings steps within the Gibbs sampler that use a Gaussian as the candidate distribution. We use the posterior mean of the MCMC samples as a point estimate of emission rate in all subsequent analysis. Section S3 in the SI contains a derivation of each conditional distribution used in the sampler.
@@ 3.3 Inference on a moving time window
The MDLQ model assumes that emission rates are constant over the timeframe of the inversion, but methane emissions on oil and gas sites are highly variable over time. We therefore propose a windowed analysis when using the MDLQ in practice. Specifically, we propose running the MDLQ model on a 30-minute inversion window that is moved forward in 10-minute increments as new data are collected. Running the MDLQ model on a shorter window would be more likely to satisfy the constant emission source assumption, but would result in higher variability in the inferred emission characteristics given the smaller sample size. We find that a 30-minute inversion length provides a compromise between these two competing effects, and we present a sensitivity study of this length in Section S4 of the SI.
@@ 3.4 Assessing the information content of each inversion window
Occasionally, there are no CMS sensors downwind of one or more sources for an entire 30-minute inversion window. When this happens, the sensors will not detect elevated concentration values if one of these sources is emitting, and as such, they provide no information about the emission state of these sources. We therefore avoid making inference on sources with no downwind sensors during a given inversion window. We can identify these sources before running the inversion via the design matrix, $X$; columns of $X$ that consist of only zero concentration predictions likely have no downwind sensor during that inversion window. In this study, we say that the “viable sources” for a given window are those whose columns of $X$ contain at least four non-zero concentration predictions. For each window, we subset $X$ to just the viable sources and run the inversion as described in Section 3.2, resulting in a rate estimate for only the viable sources. The percent of inversion windows with information on each subset of sources is given in Table 1.
@@
We note that $\boldsymbol{y}$ will be identically $\boldsymbol{0}$ for some inversion windows (e.g., from 4:00 and 8:00 in Figure 2). This corresponds to periods where the CMS sensors did not detect methane concentrations above background levels. When this occurs, we set the emission rate for all viable sources to 0 kg/hr without running the MDLQ model. As discussed above, we make no inference on sources with no downwind sensors, and as such, we do not set their emission rate to zero in this situation.
@@ 3.5 Emissions inventories and alerts
Emission mitigation using CMS is done through two primary mechanisms: emissions inventories and alerts. Here we define these two terms and describe how output from the MDLQ model can be used to create inventories and alerts to drive emission reduction.
A source-level inventory estimates how much methane a given source emitted over a certain period of time (here we use the length of the METEC study). Accurate inventories are necessary to implement regulatory penalties for excess emissions and to ensure that long term emission reduction strategies are working. We create source-level inventories by multiplying the posterior mean of $\boldsymbol{\beta}$ for each inversion window by the length of the window and then summing across the duration of the experiment (averaging the temporal domain where windows overlap). We create 95% confidence intervals on the source-level inventories by sampling from the posterior distribution of $\boldsymbol{\beta}$ for each interval, creating source-level inventories for each sample, and then selecting the $0.025^{\text{th}}$ and $0.975^{\text{th}}$ percentiles of the resulting inventory distributions. We create a site-level inventory by summing the source-level inventories. When a source has no emission rate for a given inversion window due to a lack of information (i.e., no downwind sensor), we sample from the other rate estimates for that source when calculating the inventory. This procedure is justified because there is low correlation between emission characteristics (e.g., rate) and wind direction, which determines when there are downwind sensors of the potential sources.
An emission alert contains a point estimate of which equipment are emitting at the time of the alert. Oil and gas operators use alerts to determine whether detected emissions are from normally operating equipment or from malfunctioning equipment that require immediate intervention. The proportion of posterior samples where $z_{i}=1$ gives the probability that source $i$ is emitting. We therefore create a point estimate of the emission state of source $i$ for a given inversion window as follows. If the posterior mean of $z_{i}$ is greater than the average of $z_{i}$ across all other windows, then we say that source $i$ is emitting during that inversion window. We do not create alerts when the CMS data do not contain information for a given source.
@@ 4.1 Methane emission source apportionment at METEC
We begin by showing output from the MDLQ model over an example 5-day period in Figure 5. We make two observations. First, the MDLQ model successfully distinguishes periods of emissions from periods of no emissions; it rarely infers a non-zero emission rate when no controlled releases are occurring. Second, inferred emission rates vary over the course of each release, despite the true emission rates being constant. This is because the 30-minute inversion windows are shorter than most releases, and inadequacies of the GPM will cause individual inversion windows to either over or underestimate the true emission rate.
@@
We now summarize our source apportionment results over the entire METEC experiment, focusing on the MDLQ model’s ability to produce inventories and alerts. Figure 6(a) shows site- and source-level inventories for the study period. The site-level inventory is extremely close to truth, underestimating by only 1.0%. This is promising, as site-level inventories are the primary tool for emissions accounting in US regulations (EPA, 2024), EU import standards (Council of European Union, 2023), and voluntary emission management protocols (OGMP 2.0, 2024). However, the tank inventory is overestimated by 61.4% and the east separator inventory is underestimated by -38.8%. The remaining source-level inventories are slightly underestimated with errors ranging from -3.7% to -26.3%. The cause of these source-level inventory errors is investigated fully in Section S5 of the SI, but, in short, they are likely a result of inversion windows with downwind sensors for only a subset of the window. When there is no downwind sensor for the entire window, the MDLQ model does not infer an emission rate (as discussed in Section 3.4). However, when only a subset of the window has a downwind sensor, it naively appears as if no emissions occurred during the remainder of the window and an emission rate is still inferred. This happens more often for the sources that are close to the perimeter of the site and less often for the tanks, which are centrally located. As a result of this geometry, the MDLQ model likely overestimates tank emissions to compensate for the missed emissions from the remaining sources. To test this, we keep only the inversion windows that have downwind sensors for the entire window and re-calculate the inventories. The resulting site-level inventory remains accurate (error = 1.7%) and the accuracy of the source-level inventories are much improved, with errors ranging from -23.4% to 24.2% (much closer to the inventories’ confidence intervals). This filtering, however, significantly reduces the amount of data available to apportion emissions.
@@
Figure 6: Summary of the source apportionment results for the METEC controlled releases. (a) Site- and source-level inventories created by summing the emission rate estimates across all inversion windows. Vertical lines show 95% confidence intervals. (b) Distribution of errors in the site-level emission rate estimates. (c) Alerting accuracy, where a correct localization estimate is a correct estimate of emission state (i.e., either emitting or not emitting).
@@
Figure 6(b) shows the distribution of errors in the site-level emission rate estimates across all inversion windows. The average error is -0.01 kg/hr, which makes sense given the small error in the site-level inventory from Figure 6(a). However, the inner 95% interval of the site-level errors is [-2.7, 4.0] kg/hr, meaning that there is relatively large variability in the individual site-level rate estimates. We also observed this fact in the example in Figure 5. This finding gives rise to the question: how many emission rate estimates need to be averaged before the site-level estimate is close to the truth? We investigate this question in Section S6 of the SI, and find that there is a 95% chance that the site-level error is within [-1, 1] kg/hr after averaging just 21 inversion windows (or approximately 11 hours). The fast convergence of the average is because there is low bias in the site-level rate estimates.
@@
Although this fact is not shown in Figure 6(b), we note that the coverage of the 95% credible intervals for the site-level rate estimates is only 0.58, meaning that we have missed a source of uncertainty in the MDLQ model. The missing uncertainty is very likely coming from bias in the GPM, as this model (while more accurate than comparable dispersion models) is still an oversimplification of actual atmospheric transport. We test this theory via a simulation study discussed shortly in Section 4.2.
@@
TPR 67.6% 53.9% 59.2% 64.1% 60.7% 82.7%
TNR 81.5% 80.1% 72.3% 84.1% 84.2% 76.3%
PPV 56.1% 53.3% 46.3% 60.1% 63.9% 84.8%
NPV 87.8% 80.5% 81.5% 86.3% 82.3% 73.4%
Accuracy 77.9% 72.4% 68.5% 78.7% 76.8% 80.2%
Figure 6(c) assesses the accuracy of the alerts generated by the MDLQ model. For each inversion window, we check if the localization estimate (i.e., the emission state estimate) for each source is correct. That number, ranging from 0 to 5, is plotted in Figure 6 for all inversion windows. On average, the MDLQ correctly identifies 3.79 out of the 5 possible source as either emitting or not emitting. It most frequently identifies all 5 sources correctly, doing so for 41.7% of the windows. Table 2 shows additional detection metrics at both the source- and site-level. The MDLQ model detects 82.7% of emissions at the site-level, but is only able to detect 53.9% to 67.6% of emissions at the source-level. In other words, if a piece of equipment is leaking on the site, the MDLQ model will likely detect it and can generate an alert for the site. However, determining exactly which subset of sources is emitting is a harder problem, and the MDLQ model is more prone to errors in this use case. Full source- and site-level confusion matrices are provided in Section S7 of the SI.
@@ 4.2 Simulation study
We conduct a simulation study to isolate the effect of known biases in the GPM from other potential sources of error. Specifically, we study the impact of misaligned enhancements between the GPM predictions and CMS sensor observations, examples of which can be seen in Figure 3. This behavior is a result of various inadequacies of the GPM, such as not resolving turbulent eddy motion or obstructions.
@@
$\tilde{\boldsymbol{y}}=X\boldsymbol{\beta}_{T}+\tilde{\boldsymbol{\epsilon}},$
@@
Figure 7 shows the results of this simulation study. We plot four response features for each value of $M$: the error variance, the autocorrelation coefficient, error in the emission rate estimates, and the coverage of the emission rate credible intervals. In subfigures (a)-(c), we plot the mean of each response feature across all inversion windows, with black bars showing their inner 95% range. In subfigures (c)-(d), we show results separated by emission source (in color) and when summed to the site total (in black). For subfigure (d), coverage is computed as the percent of inversion windows where the true emission rate was contained within the corresponding 95% credible interval.
Figure 7: Simulation study showing MDLQ performance under different levels of forward model bias. Red horizontal lines show the true values for each response feature. Bars show the inner 95% range for each value across the inversion windows. Colored circles correspond source-level results, and black squares correspond to site-level results.
We make three observations. First, all parameters are successfully estimated in the absence of forward model bias, as expected, with source-level coverages close to 95%. The site-level coverage is lower than 95% because the site-level credible interval will not contain the truth when any of the source-level credible intervals do not contain the truth. Since errors in the source-level rate estimates do not always occur during the same inversion window, the site-level credible interval will contain the truth for less than the expected 95% of inversion windows. Second, we see a diminishing ability to estimate $\sigma^{2}$ and $r$ as the degree of forward model bias increases, but emission rate estimates remain largely unbiased with a slight tendency to underestimate. This is promising, as $\boldsymbol{\beta}$ are the parameters of direct interest for methane emission source apportionment. The overestimation of $\sigma^{2}$ is because the misaligned concentration enhancements result in extremely large errors, either positive or negative, which the Gaussian error model is not flexible enough to accommodate. Some of these errors are likely misinterpreted as autocorrelation, resulting in overestimates of $r$. Finally, source and site-level coverages drop as the degree of forward model bias increases and align with the coverage values seen during the METEC experiment. A drop in coverage is expected, as the misaligned enhancements result in either over or underestimation of $\boldsymbol{\beta}$, such that many of the 95% credible intervals (which are not aware of the GPM bias) no longer contain the truth. Site-level coverage for the METEC experiment was 0.58, and source-level coverages ranged from 0.68 to 0.76. These values align closely with the coverages seen in the simulation study, providing evidence that GPM bias is indeed the cause of the coverage issues during the METEC controlled releases.
Section S8 in the SI contains an additional simulation study demonstrating the importance of modeling autocorrelation in $\boldsymbol{\epsilon}$; failing to do so results in overestimated $\boldsymbol{\beta}$ and low coverages.

### v2

@@ 3.1 Bayesian hierarchical model
$\theta_{i}$ $\sim\text{Beta}(a_{i},b_{i})$ (9)
$\tau_{i}^{2}$ $\sim\text{Inv-Gamma}(c_{i},d_{i})$ (10)
$\sigma^{2}$ $\sim\text{Inv-Gamma}(\nu/2,\nu/2)$ (11)
$\nu$ $\sim\text{Inv-Gamma}(\alpha_{1},\alpha_{2})$ (12)
$r$ $\sim\text{Uniform}(0,1).$ (13)
Fixed hyperparameters $a_{i}$, $b_{i}$, $c_{i}$, and $d_{i}$ control any prior information about the probability that $\beta_{i}$ is non-zero and the scale of $\beta_{i}$. For example, heaters on oil and gas sites are not used in the summer, and as such the $a_{i}$ and $b_{i}$ hyperparameters could be set such that $\theta_{i}$ has more mass close to zero in the prior during these months. In this article, however, we use a uniform prior for $\theta_{i}$ by setting $a_{i}=b_{i}=1$ and relatively uninformative priors for the remaining parameters by setting $c_{i}=d_{i}=\alpha_{1}=\alpha_{2}=1$. These values were selected to reflect prior knowledge that oil and gas methane emissions (on all sites, not just METEC) are mostly concentrated at small values between 0 and 10 kg/hr, with rare emissions occurring on the order of 100 to 100,000 kg/hr. In particular, setting $c_{i}=d_{i}=1$ results in an Inverse Gamma prior for $\tau^{2}_{i}$ with undefined mean and variance; this allows for very large $\tau^{2}_{i}$ to be sampled without putting too much mass at extreme, unrealistic values. Emissions estimates from the MDLQ model are not highly sensitive to the choice of $c_{i}$ and $d_{i}$ for similarly realistic values. See SI Section S12 for more details and a sensitivity study of these prior hyperparameters.
The prior for $\sigma^{2}$ was set such that, in the prior model, the $\boldsymbol{\epsilon}$ are distributed according to a Student’s t-distribution with $\nu$ degrees of freedom, as we discovered that $\boldsymbol{\epsilon}$ can be heavier tailed than a Gaussian through basic model diagnostics. This prior belief can be updated by the data. We find that emission rate estimates from the MDLQ model are not sensitive to the choice of $\alpha_{1}$ and $\alpha_{2}$; see SI Section S12 for details.
@@ 3.2 Posterior inference
We sample from the joint posterior of all model parameters using a Markov chain Monte Carlo (MCMC) algorithm. We describe our MCMC algorithm below, but see Gelman et al. (2015) for a more complete discussion of Bayesian inference via MCMC. While samples are obtained for all model parameters, the emission rates, $\boldsymbol{\beta}$, and indicator variables, $\boldsymbol{z}$, are the most useful parameters for methane emission source apportionment on oil and gas sites.
We use a Metropolis-within-Gibbs algorithm to sample from the posterior. We are able to derive a closed form expression for the conditional distributions of all model parameters except $\nu$ and $r$, resulting in a fast sampler that mixes well. For $\nu$ and $r$, we update using two Metropolis-Hastings steps within the Gibbs sampler that use a Gaussian as the candidate distribution. We use the posterior mean of the MCMC samples as a point estimate of emission rate in all subsequent analysis. Section S3 in the SI contains a derivation of each conditional distribution used in the sampler.
@@ 3.3 Inference on a moving time window
The MDLQ model assumes that emission rates are constant over the temporal domain of the inversion, $\mathcal{D}=\{1,...,l\}$ even though methane emissions on oil and gas sites are known to vary over time. As such, this assumption breaks down for large $l$. To address this issue, we partition the full time series of concentration observations into consecutive, non-overlapping windows of length $l=30$ minutes and treat each window as an independent inverse problem. For a given window, $\boldsymbol{y}\in\mathbb{R}^{n}$ denotes the concatenated minute-frequency concentration observations across all sensors within that 30-minute period, exactly as defined in Section 3.1, with $n=l\times m$. Likewise, the corresponding design matrix $X\in\mathbb{R}^{n\times p}$ is constructed from the GPM simulations for the same 30-minute period. For the METEC experiment, with $l=30$, $m=10$, and $p=5$, we get $\boldsymbol{y}\in\mathbb{R}^{300}$ and $X\in\mathbb{R}^{300\times 5}$ for each window.
This windowed approach yields a piecewise-constant representation of emission behavior over time, with temporal resolution determined by the window length, $l$. The choice of a 30-minute window reflects a tradeoff between model assumptions and stability of the inferred parameters. Shorter windows are more likely to satisfy the constant emission assumption but provide fewer observations, leading to higher uncertainty and increased sensitivity to transport model errors. Longer windows provide more data for each inversion but are increasingly likely to violate the constant emission rate assumption. Based on exploratory analysis and a formal sensitivity study presented in Section S4 of the SI, we find that a 30-minute window provides a practical compromise for the METEC dataset.
@@ 3.4 Assessing the information content of each inversion window
Occasionally, there are no CMS sensors downwind of one or more sources for an entire 30-minute inversion window. When this happens, the sensors will not detect elevated concentration values if one of these sources is emitting, and as such, they provide no information about the emission state of these sources. We therefore avoid making inference on sources with no downwind sensors during a given inversion window. We can identify these sources before running the inversion via the design matrix, $X$; columns of $X$ that consist of only zero concentration predictions likely have no downwind sensor during that inversion window.
In this study, we say that the “viable sources” for a given window are those whose columns of $X$ contain at least four non-zero concentration predictions. We intentionally chose a low threshold so that weakly informative inversion windows are still included in our analysis. This allows us to gauge how incomplete sensor coverage influences source apportionment accuracy and is discussed further in Section 4.1 and SI Section S5. For each window, we subset $X$ and $\boldsymbol{\beta}$ to just the viable sources and run the inversion as described in Section 3.2, resulting in a rate estimate for only the viable sources. The percent of inversion windows with information on each subset of sources is given in Table 1.
@@
We note that $\boldsymbol{y}$ will be identically $\boldsymbol{0}$ for some inversion windows (e.g., from 4:00 to 8:00 in Figure 2). This corresponds to periods where the CMS sensors did not detect methane concentrations above background levels. When this occurs, we set the emission rate for all viable sources to 0 kg/hr without running the MDLQ model. As discussed above, we make no inference on sources with no downwind sensors, and as such, we do not set their emission rate to zero in this situation.
@@ 3.5 Emissions inventories and alerts
Emission mitigation using CMS is done through two primary mechanisms: emissions inventories and alerts. Below, we define these terms and describe how they are constructed using MDLQ output.
A source-level inventory estimates the total emitted methane from a given source over a specific period of time (here, the duration of the METEC study). Inventories are used to implement regulatory penalties for excess emissions and to evaluate long-term emission reduction strategies. We construct source-level inventories by multiplying the posterior mean of $\boldsymbol{\beta}$ in each inversion window by the length of the window and then summing across the experiment. We quantify uncertainty by sampling from the posterior of $\boldsymbol{\beta}$ in each window, constructing a source-level inventory for each sample, and reporting the $0.025^{\text{th}}$ and $0.975^{\text{th}}$ percentiles as a 95% credible interval on the inventory. We construct a site-level inventory by summing the source-level inventories. For windows where a source has no emissions estimate due to a lack of information (i.e., no downwind sensor), we sample from that source’s emissions estimates in other windows. This approach is justified by the low correlation between emission rate and wind direction, which determines when a source has downwind sensors.
Emission alerts provide a point estimate of which sources are emitting during a given inversion window. They are used by oil and gas operators to determine whether detected emissions are from normally operating equipment or from malfunctioning equipment that require immediate intervention. The posterior probability that source $i$ is emitting is given by the proportion of samples with $z_{i}=1$. We therefore classify source $i$ as emitting in a given window if the posterior mean of $z_{i}$ exceeds its average across all other windows. We do not create alerts when the CMS data do not contain information for a given source.
@@ 4.1 Methane emission source apportionment at METEC
We begin by showing output from the MDLQ model over an example 5-day period in Figure 5. We make two observations. First, the MDLQ model successfully distinguishes periods of emissions from periods of no emissions; it rarely infers a non-zero emission rate when no controlled releases are occurring, and vice versa (see Table 2 for details). Second, inferred emission rates vary over the course of each release, despite the true emission rates being constant. This is because the 30-minute inversion windows are shorter than most releases, and inadequacies of the GPM will often cause individual inversion windows to either over or underestimate the true emission rate. We discuss MDLQ model refinements that may mitigate this behavior in Section 5.
@@
We now summarize our source apportionment results over the entire METEC experiment, focusing on the MDLQ model’s ability to produce inventories and alerts. Figure 6(a) shows site- and source-level inventories for the study period. The site-level inventory is close to the truth, overestimating by only 4.5%. This is promising, as site-level inventories are the primary tool for emissions accounting in US regulations (EPA, 2024), EU import standards (Council of European Union, 2023), and voluntary emission management protocols (OGMP 2.0, 2024). However, the tank inventory is overestimated by 81.8% and the east separator inventory is underestimated by -37.7%. The remaining source-level inventories are slightly underestimated with errors ranging from -8.9% to -22.9%. The cause of these source-level inventory errors is investigated fully in Section S5 of the SI, but in short, they are likely a result of inversion windows with downwind sensors for only a subset of the window. When there is no downwind sensor for the entire window, the MDLQ model does not infer an emission rate (as discussed in Section 3.4). However, when only a subset of the window has a downwind sensor, it naively appears as if no emissions occurred during the remainder of the window and an emission rate is still inferred. This happens more often for the sources that are close to the perimeter of the site and less often for the tanks, which are centrally located. As a result of this geometry, the MDLQ model likely overestimates tank emissions to compensate for the missed emissions from the remaining sources. To test this, we keep only the inversion windows that have downwind sensors for the entire window and re-calculate the inventories. The resulting site-level inventory remains accurate (error = 2.0%) and the accuracy of the source-level inventories are much improved, with errors ranging from -22.0% to 20.4% (much closer to the inventories’ credible intervals). This filtering, however, significantly reduces the amount of data available to apportion emissions.
@@
Figure 6: Summary of the source apportionment results for the METEC controlled releases. (a) Site- and source-level inventories created by summing the emission rate estimates across all inversion windows. Vertical lines show 95% credible intervals. (b) Distribution of errors in the site-level emission rate estimates. (c) Alerting accuracy, where a correct localization estimate is a correct estimate of emission state (i.e., either emitting or not emitting).
@@
Figure 6(b) shows the distribution of errors in the site-level emission rate estimates across all inversion windows. The average error is 0.06 kg/hr, which makes sense given the small error in the site-level inventory from Figure 6(a). However, the inner 95% interval of the site-level errors is [-2.7, 4.5] kg/hr, meaning that there is relatively large variability in the individual site-level rate estimates. We also observed this fact in the example in Figure 5. This finding gives rise to the question: how many emission rate estimates need to be averaged before the site-level estimate is close to the truth? We investigate this question by resampling from the site-level error distribution at different sample sizes (see SI Section S6); we find that 95% of the site-level sample means are within [-1, 1] kg/hr after averaging just 34 inversion windows (or approximately 17 hours). The average emissions estimate converges quickly because there are no large outliers in the site-level error distribution for the METEC experiment.
@@
Although this fact is not shown in Figure 6(b), we note that the coverage of the 95% credible intervals for the site-level rate estimates is only 0.66, meaning that we have missed a source of uncertainty in the MDLQ model. The missing uncertainty is very likely coming from bias in the GPM, as this model (while more accurate than comparable dispersion models) is still an oversimplification of actual atmospheric transport. We test this theory via a simulation study discussed shortly in Section 4.2.
Figure 6(c) assesses the accuracy of the alerts generated by the MDLQ model. For each inversion window, we check if the localization estimate (i.e., the emission state estimate) for each source is correct. That number, ranging from 0 to 5, is plotted in Figure 6 for all inversion windows. On average, the MDLQ correctly identifies 4.01 out of the 5 possible source as either emitting or not emitting. It most frequently identifies all 5 sources correctly, doing so for 46.2% of the windows. Table 2 shows additional detection metrics at both the source- and site-level. The MDLQ model detects 83.8% of emissions at the site-level, but is only able to detect 56.2% to 67.9% of emissions at the source-level. In other words, if a piece of equipment is emitting, the MDLQ model will likely detect it and generate an alert for the site. However, determining exactly which subset of sources is emitting is a harder problem, and the MDLQ model is more prone to errors in this use case. Full source- and site-level confusion matrices are provided in Section S7 of the SI.
@@
TPR 67.9% 56.2% 60.6% 64.4% 61.2% 83.8%
TNR 90.5% 90.4% 77.4% 90.7% 92.3% 91.2%
PPV 73.7% 73.6% 55.0% 74.5% 80.6% 95.1%
NPV 87.8% 81.3% 81.1% 85.8% 82.0% 73.4%
Accuracy 84.1% 79.4% 72.1% 82.9% 81.7% 86.2%
@@ 4.2 Simulation study
We conduct a simulation study to isolate the effect of known biases in the GPM from other potential sources of error. Specifically, we study the impact of misaligned concentration enhancements between the GPM predictions and CMS sensor observations, examples of which can be seen in Figure 3. This behavior is a result of various inadequacies of the GPM, such as not resolving turbulent eddy motion or obstructions.
Figure 7: Simulation study showing MDLQ performance under different levels of forward model bias. Red horizontal lines show the true values for each response feature. Bars show the inner 95% range for each value across the inversion windows. Colored shapes correspond to source-level results, and black circles correspond to site-level results.
@@
$\tilde{\boldsymbol{y}}=X\boldsymbol{\beta}_{T}+\tilde{\boldsymbol{\epsilon}}$, (14)
@@
Figure 7 shows the results of this simulation study. We plot four response features for each value of $M$: the marginal error variance, the autocorrelation coefficient, error in the emission rate estimates, and the coverage of the emission rate credible intervals. In subfigures (a)-(c), we plot the mean of each response feature across all inversion windows, with black bars showing their inner 95% range. In subfigures (c)-(d), we show results separated by emission source and when summed to the site-level. For subfigure (d), coverage is computed as the percent of inversion windows where the true emission rate was contained within the corresponding 95% credible interval.
We make three observations. First, all parameters are accurately estimated in the absence of forward model bias, as expected, with source-level coverages close to 95%. Second, we see a diminishing ability to estimate $\sigma^{2}$ and $r$ as the degree of forward model bias increases, but emission rate estimates remain largely unbiased with a slight tendency to underestimate. This is promising, as $\boldsymbol{\beta}$ are the parameters of direct interest for methane emission source apportionment. The overestimation of $\sigma^{2}$ is because the misaligned concentration enhancements result in extremely large errors, either positive or negative, which the Gaussian error model is not flexible enough to accommodate. Some of these errors are likely misinterpreted as autocorrelation, resulting in overestimates of $r$. Finally, source and site-level coverages drop as the degree of forward model bias increases and approach the coverage from the METEC experiment. A drop in coverage is expected, as the misaligned enhancements result in either over or underestimation of $\boldsymbol{\beta}$, such that many of the 95% credible intervals (which are not aware of the GPM bias) no longer contain the truth. Site-level coverage for the METEC experiment was 0.66, and source-level coverages ranged from 0.71 to 0.81. The fact that coverage values from the simulation study approach those from the METEC experiment provides evidence that GPM bias is indeed a large cause of the underestimated uncertainty in the MDLQ model. The fact that the simulation study coverages are not identical to the real data coverages means that there are additional missing sources of uncertainty beyond misaligned concentration enhancements.
Section S8 in the SI contains another simulation study demonstrating the importance of modeling autocorrelation in $\boldsymbol{\epsilon}$; failing to do so results in overestimated $\boldsymbol{\beta}$ and low coverages.